# What Makes for Good Image Captions?

### v2

@@ 1 Introduction
Image captioning, the process of translating visual content into natural language descriptions, serving pivotal roles in real-world applications ranging from assisting visually impaired individuals [1, 2, 3, 4] to facilitating content-based image retrieval [5, 6, 7, 8, 9, 10]. Over the last decade, the field of image captioning has witnessed substantial progress, primarily driven by advancements in deep neural nets and the availability of large-scale high-quality image-text datasets.
Despite empirical advancements, several fundamental questions remain unanswered: What makes for good image captions? Which properties should they possess, and how can we measure them? Some existing models can generate captions closely resembling single-sentence human annotations [11], but these may not be adequate for use cases where more comprehensive coverage of fine-grained visual information is required. Conversely, recent Large Vision Language Models (LVLMs) [12, 13] have demonstrated the ability to generate multi-paragraph detailed image descriptions [14, 15, 12, 16]. Yet, longer captions can sometimes be less accurate, hallucinate content, or put excessive emphasis on irrelevant details while omitting important ones.
Recognizing the absence of a universal standard for ideal captions, this work aims to establish well-defined principles for image captioning that address varying task requirements. We introduce an information-theoretic framework based on semantic communication [17, 18] and the information bottleneck principle [19, 20, 21]. By leveraging this perspective, we formulate an objective function for image captioning that strikes a balance among three key criteria:
@@ 2 Proposed Framework
In this section, we provide a theoretical framework for image captioning as depicted in Figure 1. First, we formulate the task of image captioning by applying the concept of semantic units [17, 18]. We suppose that an image is an observation of a latent variable in a semantic space characterized by semantic units. An image captioning model will generate a caption for the image, and the caption can be mapped back to the latent semantic space and compared with the source semantic latent variable.
Based on this framework, we then introduce our proposed objectives inspired by the information bottleneck principle [19, 20, 21] for feature representation learning [21]. In our framework, we consider that the overall image captioning objective is composed of a “information sufficiency” term, a “minimal redundancy” term, and a “human comprehensibility” term.
Figure 1: Overview of our formulation. Some latent variable $X$ in a latent semantic space $\mathcal{S}$ generates image $\tilde{X}$ in data space $\mathcal{D}$. The image $\tilde{X}$ is then captioned by $f_{\theta}$ producing a caption $\tilde{Y}$ which can be mapped back to the original latent space as $Y$. The semantic-level error $Z=Y-X$ measures the difference between the source semantics $X$ and received semantics $Y$.
@@ 2.1 Formulation
We assume that meaning arises from a set of independent and discrete units called semantic units in a semantic space, and images and captions are observations of some latent variables in this semantic space [17, 18]. Following [17, 18], we define a semantic unit and the semantic space as:
Definition 1 (Semantic Units and Semantic Space)
A semantic unit represents an atomic piece of information. The set of all possible semantic units is denoted by $\Omega=\{w_{i}\}_{i=1}^{n}$. A semantic space is an $n$-dimensional space $\mathcal{S}\in[0,1]^{n}$ where the value of the $i$-th dimension of represents the probability of the presence of the corresponding semantic unit $p(w_{i})$.
The $\Omega$ encompasses a wide range of semantic units, similar to how a vocabulary contains diverse words. Each point in $\mathcal{S}$ corresponds to a specific combination of semantic units, corresponding to different meanings. Adopting the semantic space and its probabilistic interpretation in [17, 22, 23], we can apply the classical information theory [24] to operate at the semantic level, and make information in images and captions to be comparable.
@@
Definition 2 (Image Captioning)
An image captioning model $f$ parameterized by $\theta$ operates in the observed data spaces $f_{\theta}:\mathcal{D_{\mathrm{image}}}\rightarrow\mathcal{D_{\mathrm{caption}}}$, it translates an image into a caption, i.e., $\tilde{Y}=f_{\theta}(\tilde{X})$. $\tilde{X}$ is generated from some source latent variable $\tilde{X}=g(X)$ while $\tilde{Y}$ can be converted back to the latent semantic space $Y=h(\tilde{Y})$. Let $Z=Y-X\in[-1,1]^{n}$ denote the error between the source and recovered semantics caused by parameters $\theta$.
@@ 2.2 Objectives
The image captioning process can be compared to a communication system [24] where information source $X$ is converted to signal $\tilde{X}$ by a lossless source encoder $g$, transmitted through a noisy channel $f_{\theta}$, and the received signal $\tilde{Y}$ is losslessly decoded by $h$, giving the final received information $Y$. From this communication system perspective, one might say that the optimal $\theta^{*}$ could just be the one that minimizes the error $||Z||$. However, this requirement is unrealistic as it would result in extremely long captions that losslessly encode both high-level semantic information and all the low-level irrelevant information.
Therefore, we apply the information bottleneck principle [19, 20, 21] to evaluate this system. Information bottleneck is a generalization of rate-distortion theory for lossy data compression, it requires a representation (which in our case is $\tilde{Y}$) to have maximal mutual information with some information $T$ that is required fulfill the task requirements (i.e., high information sufficiency), while having minimal mutual information with the input $X$ (minimal redundancy). The desired minimal sufficient representation can be given as $\tilde{Y}^{*}=\arg\max I(\tilde{Y};T)-\beta I(X;\tilde{Y})$ where $\beta$ is a Lagrange multiplier. If the captioning model is deterministic (given $X$, it always produces the same $\tilde{Y}$) then we have $H(\tilde{Y}|X)=0$. Since $I(X;\tilde{Y})=H(\tilde{Y})-H(\tilde{Y}|X)$, the minimal sufficient representation can be written as:
$\tilde{Y}^{*}=\arg\underset{\tilde{Y}}{\max}\ I(\tilde{Y};T)-\beta H(\tilde{Y}).$ (1)
@@
Next, we give a formal definition of the information sufficiency objective of image captioning with importance of semantic units [17].
Definition 3 (Information Sufficiency)
For given $X$, let a latent variable $T\in\mathcal{S}$ represent the task-relevant information in $X$, and let an importance variable $A\in[0,1]^{n}$ denote the importance scores of different semantic units. The $A$ is derived from $X$ by an underlying mapping, thus being dependent on $X$. The $T$ is produced by a point-wise product between $A$ and $X$, thus $T=A\odot X$. For generated image captions $\tilde{Y}=f_{\theta}(\tilde{X})$, the information sufficiency objective is:
@@
In the importance variable, $A_{i}=1$ means the semantic units $w_{i}$ are very important in the image, while $A_{i}=0$ means $w_{i}$ is irrelevant. It behaves similarly to the attention mechanism [25], which also produces a heatmap between zero to one according to the given input. Note that here the $A$ is not binary, the continual nature of $A$ gives a good property to $J_{\mathrm{suf}}(\theta)$: when the “budget” is limited (as there are also other objectives to optimize), more semantic units with higher importance score will be retained while less important ones will be discarded.
Definition 4 (Minimal Redundancy)
The minimal redundancy objective encourages the image captioning model $f_{\theta}$ to eliminate irrelevant information, it is given by measuring the entropy of generated captions:
$J_{\mathrm{min}}(\theta)=-H(\tilde{Y}).$ (3)
@@
Definition 5 (Human Comprehensibility)
Let $P_{\tilde{Y}}$ denote the probabilistic distribution of model-generated captions over $\mathcal{D_{\mathrm{caption}}}$, and let $P_{\mathrm{lang}}$ denote the distribution of human interpretable natural language. Given a certain statistical divergence measurement $D$, the human comprehensibility objective is:
@@
3 Example Application of the Framework
3.1 Method: The Pyramid of Captions
In this section, we introduce the Pyramid of Captions (PoCa) method, which showcases the applicability of our framework to image captionoing research. The key intuition behind the PoCa method is that we can have a more accurate and detailed caption by ensembling multiple captions.We propose to split an image into multiple local patches generate local captions for each patch, and fuse the captions to obtain a higher-quality caption for the global image.
@@ 3.1 Method: The Pyramid of Captions
We apply an image captioning model $f_{\theta}$ to the local patches and obtain a set of local captions: $\{\tilde{Y}^{[j]}=f_{\theta}(\tilde{X}^{[j]})\}_{j=1}^{m}$, and also generate a caption for the global image: $\tilde{Y}=f_{\theta}(\tilde{X})$. The local and global information will be fused by a merging function $\sigma_{\mathrm{merge}}$ operating in $\mathcal{D}_{\mathrm{caption}}$:
@@
We adopt text-only LLMs as the merging function $\sigma_{\mathrm{merge}}$. Table 2 provides an example of merging using an LLM, where we instruct it to generate a merged caption that incorporates both local and global information.
@@
3.2 PoCa Gives Better Captions (Provably)
@@
Assumption 1 (Uncertainty-aware content-dependent error)
The error $Z$ produced by the image captioning model $f_{\theta}$ is dependent on the input semantics $X$. Therefore, it can be expressed as a deterministic function of $X$:
$||Z_{i}||=\varphi(X_{i}),$ (8)
@@
Assumption 2 (Local-global relationship of image semantics)
The $\sigma_{\mathrm{split}}$ splits an image into local patches. The latent semantic variables corresponding to local patches satisfy the following relationship with global semantics:
@@
Assumption 3 (Local-global aggregation of caption semantics)
The function $\sigma_{\mathrm{merge}}$ merges the global and global captions. The latent semantic variable corresponding to the merged caption is a weighted combination of the global and local semantics:
@@
Theorem 1 (PoCa method reduces semantic error)
@@
Under Assumptions 1-3, the PoCa method is guaranteed to have the smaller error $Z_{\mathrm{merged}}$ than $Z$, i.e.,
$||Z_{\mathrm{merged}}||\leq||Z||.$ (11)
@@
Since $A$ is non-negative, Theorem 1 implies non-decreasing information sufficiency; if merging does not increase redundancy or decrease interpretability, then the overall quality of caption becomes better. This also aligns with the findings in [26] that smaller-scale models combined can be as effective as a larger-scale model. Additionally, it is worth noting that our assumptions require linear combinations of semantics, while it may not hold in practice for images with more complex structures of semantics.
3.3 PoCa Gives Better Captions (Empirically)
We conduct quantitative evaluations to study whether the PoCa method can improve the caption quality. We adopt the VQA-v2 [27] dataset, which is built upon the MS-COCO [11] dataset and contains multiple questions per image. The questions serve as a proxy for the importance score $A$ in Definition 3, and the accuracy of the text-only LLM (LLaMA2-Chat-13B, prompt shown in Appendix Table 5) generated answers becomes an estimation of the information sufficiency term. See Appendix for more implementation details.
Figure 2: Evaluation of Sufficiency. VQA accuracy using captions generated by different image captioning models and the proposed PoCa method. The PoCa method consistently improves the VQA performance across all models and caption types.
Figure 2 provides the evaluation results on 5,000 questions in the VQA-v2 validation split. As can be seen, our proposed PoCa method (green) consistently yields performance gains across all three examined LVLMs. The scale of improvement ranges from 0.48% (MobileVLM-v2 detailed captions) to 2.10% (LLaVA-1.5 short captions). Interestingly, we find that detailed captions do not necessarily correlate with better information coverage, as the detailed captions generated by MobileVLM-v2 underperform the single-sentence captions generated by InternVL. The comparison also shows that human annotations may not be optimal for certain scenarios, since several groups of LVLM-generated captions can yield higher VQA accuracy compared to that of human annotators.
It is crucial to evaluate whether the performance gain brought by PoCa is achieved by significantly sacrificing other objectives. In Table 1, we present the length statistics. We calculate the average number of words in default captions and PoCa captions and note their differences in the “±$\pm\Delta$” column. The results show that PoCa does not exhibit a significant trend of either increasing or decreasing the length of captions. Among the six comparisons, PoCa compresses the length in four cases and extends the length in two cases. This empirically demonstrates that using LLMs as $\sigma_{\mathrm{merge}}$ in the PoCa model does not significantly violate the minimal redundancy objective $H(\tilde{Y}_{\mathrm{merged}})$.
Table 1: Evaluation of Redundancy. Caption lengths statics between default and PoCa captions.
LVLM VQAv2 [27] Img2P [28]
Default PoCA ±$\pm\Delta$ Default PoCA ±$\pm\Delta$
MobileVLM-v2-1.7B 54.1 78.2 +24.1 61.6 47.0 -14.6
LLaVA-1.5-7B 82.7 74.7 -8.0 93.2 133.4 +40.2
InternVL 158.3 93.4 -65.0 177.4 176.2 -1.2
4 Conclusion
Appendix
Appendix A Related Work
A.1 Image Captioning
Image captioning lies at the intersection of computer vision and natural language processing, requiring both accurate visual recognition and coherent language generation abilities. In Section 2 we formally defined the objectives for this task, but it is worth noting that current efforts do not explicitly optimize that objective. The primary challenge is the difficulty of back-propagation through the discrete textual space $\mathcal{D}_{\mathrm{caption}}$, and efforts addressing this challenge involve adopting reinforcement learning [29, 30] or aligning continuous latent spaces with language spaces [31, 32]. However, these approaches suffer from training instability and less satisfactory language coherence.
Most current methods rely on a surrogate methodology, where they use human annotators $f_{\mathrm{human}}$ to write captions and train models to imitate those captions. The underlying assumption is that human-written captions optimize the objective $J(\mathrm{human})$, which is achieved by providing instructions to crowd-sourced caption annotators. For example, the instructions for MS COCO Caption annotation [11] include “Describe all the important parts of the scene” and “Do not describe unimportant details”, which are respectively connected to the information sufficiency term and minimal redundancy term in our objective.
An important trend in the image captioning field is the increasing focus on the comprehensiveness of image captions. As mentioned earlier, this represents a decreased length penalty (smaller weight $\beta$ for the minimal redundancy term) and more emphasis on the information sufficiency term. In recent years, there has been an increasing number of high-quality detailed captioning datasets for this target, such as human-annotated image paragraph captioning [28], Densely Captioned Images (DCI) [14], and pseudo-labeled datasets, including LLaVA-Detailed-Captions [12], ShareGPT4V [33], AS-1B [16], etc. However, detailed caption annotation is much more expensive than previous single-sentence annotation, while automated caption labeling exhibits a high risk of hallucination.
A.2 Vision-Language Learning in the Era of Large Language Models
Various methods have been explored for enabling vision-language learning in LLMs. One line of work focusing on vision-language alignment during pretraining [34, 35, 36, 37, 38, 39, 40, 41, 42, 43], allowing the model to jointly learn a shared vision-language latent space. The other line of work, improve the vision-language training efficiency by aligning the vision representation into the language space of LLMs by only training the visual encoder module or a vision-language projection matrix [44, 45, 12, 46, 47, 48]. These two lines of works enable vision-language alignment, enabling various joint vision and language modalities prompting methods such as robot manipulation prompting [49, 50] and multimodal in-context learning [51, 52].
Unlike the other two directions, another line of work exploits the reasoning and planning ability of LLMs allowing zero-shot multimodal vision-language inference by extracting the information from the vision modality into a textual description and performing inference through frozen LLMs [53, 54, 55, 56]. Recent works in this direction showcase remarkable VQA ability through answer heuristics generation [57] and enabling object tagging and image editing through visual programming [58]. Inspired by this line of work, our work introduces a zero-shot hierarchical image captioning approach that relies on the reasoning ability of LLMs to aggregate information from local and global captions.
Appendix B Proof for Theorem 1
Theorem 1 (PoCa method reduces semantic error)
Under Assumptions 1-3, the PoCa method is guaranteed to have a smaller error $Z_{\mathrm{merged}}=Y_{\mathrm{merged}}-X$ than $Z$, i.e.,
$||Z_{\mathrm{merged}}||\leq||Z||.$ (12)
@@ Appendix B Proof for Theorem 1
$\Delta_{\mathrm{PoC},i}$ $=||Z_{i}||-||Z_{\mathrm{merged},i}||$ (17)
$=||Z_{i}||-||\eta Z_{i}+(1-\eta)\sum_{j}^{m}\alpha_{j}Z^{[j]}_{i}||$ (18)
$\geq||Z_{i}||-(\eta||Z_{i}||+(1-\eta)\sum_{j}^{m}\alpha_{j}||Z^{[j]}_{i}||)$ (19)
$=(1-\eta)(||Z_{i}||-\sum_{j}^{m}\alpha_{j}||Z^{[j]}_{i}||)$ (20)
$=(1-\eta)(\varphi(X_{i})-\sum_{j}^{m}\alpha_{j}\varphi(X^{[j]}_{i})).$ (21)
To yield (19) from (18), we apply the triangle inequality, which states that for any two vectors $a$ and $b$, $||a+b||\leq||a||+||b||$. Finally, we apply Assumption 1 to obtain (21), which states $||Z_{i}||=\varphi(X_{i})$ and $||Z^{[j]}_{i}||=\varphi(X^{[j]}_{i})$.
@@
The first inequality follows directly from the lower bound for $\Delta_{\mathrm{PoC},i}$ derived earlier. The second inequality follows from the Jensen’s inequality result, which states that $\varphi(X_{i})\geq\sum_{j}^{m}\alpha_{j}\varphi(X^{[j]}_{i})$. Since $01-\eta>$ 0 (as $\eta\in(0,1)$ by Assumption 3), the product of $(1-\eta)$ and a non-negative term $(\varphi(X_{i})-\sum_{j}^{m}\alpha_{j}\varphi(X^{[j]}_{i}))$ should also be non-negative, thus proving that $\Delta_{\mathrm{PoC},i}\geq 0$ for all $i$. Therefore, we have:
$||Z_{\mathrm{merged}}||\leq||Z||.$ (24)
@@
Appendix C Implementation Details
C.1 Image Captioning Models
We employ three groups of Large Vision Language Models (LVLMs) as the image captioning models: LLaVA-1.5 [59], MobileVLM v2 [60], and InternVL [61], among them:
  • LLaVA-1.5 series is a popular LVLM with two variants: LLaVA-1.5-7B and LLaVA-1.5-13B, which adopt Vicuna-7B and Vicuna-13B as their Language Models (LLMs), respectively.
  • MobileVLM v2 is a family of efficient LVLMs with smaller scales, and we utilize its MobileVLM-v2-1.7B and MobileVLM-v2-3B models.
  • InternVL is one of the top-performing publicly available LVLMs. We use its InternVL-Chat-Chinese-V1-2-Plus model, which is based on the Yi-34B LLM and has a total of 40.1B parameters.

### v3

@@ Abstract
What Makes for Good Image Captions?
Delong Chen1,2 Samuel Cahyawijaya1 Etsuko Ishii1 Ho Shu Chan1Yejin Bang 1,2Pascale Fung 1,2 1HKUST 2Meta FAIR Paris
@@ 1 Introduction
Image captioning, the process of translating visual content into natural language descriptions, serving pivotal roles in real-world applications ranging from assisting visually impaired individuals Gurari et al. (2020); Rane et al. (2021); Guo et al. (2022); Ganesan et al. (2022) to facilitating content-based image retrieval Gudivada and Raghavan (1995); Srihari (1995); Datta et al. (2005); da Silva Torres and Falcao (2006); Jain et al. (2015); Li et al. (2021). Over the last decade, the field of image captioning has witnessed substantial progress, primarily driven by advancements in deep neural nets and the availability of large-scale high-quality image-text datasets.
Despite empirical advancements, several fundamental questions remain unanswered: What makes for good image captions? Which properties should they possess, and how can we measure them? Some existing models can generate captions closely resembling single-sentence human annotations Chen et al. (2015), but these may not be adequate for use cases where more comprehensive coverage of fine-grained visual information is required. Conversely, recent Large Vision Language Models (LVLMs) Liu et al. (2023c); Dai et al. (2024) have demonstrated the ability to generate multi-paragraph detailed image descriptions Urbanek et al. (2023); Cho et al. (2023); Liu et al. (2023c); Wang et al. (2024). Yet, longer captions can sometimes be less accurate, hallucinate content, or put excessive emphasis on irrelevant details while omitting important ones.
Recognizing the absence of a universal standard for ideal captions, this work aims to establish well-defined principles for image captioning that address varying task requirements. We introduce an information-theoretic framework based on semantic communication Peyrard (2019); Zhong (2017) and the information bottleneck principle Tishby et al. (2000); Shwartz-Ziv and Tishby (2017); Tsai et al. (2021). By leveraging this perspective, we formulate an objective function for image captioning that strikes a balance among three key criteria:
@@ 2 Proposed Framework
In this section, we provide a theoretical framework for image captioning as depicted in Figure 1. First, we formulate the task of image captioning by applying the concept of semantic units Peyrard (2019); Zhong (2017). We suppose that an image is an observation of a latent variable in a semantic space characterized by semantic units. An image captioning model will generate a caption for the image, and the caption can be mapped back to the latent semantic space and compared with the source semantic latent variable.
Based on this framework, we then introduce our proposed objectives inspired by the information bottleneck principle Tishby et al. (2000); Shwartz-Ziv and Tishby (2017); Tsai et al. (2021) for feature representation learning Tsai et al. (2021). In our framework, we consider that the overall image captioning objective is composed of a “information sufficiency” term, a “minimal redundancy” term, and a “human comprehensibility” term.
Figure 1: Overview of our formulation. Some latent variable $X$ in a latent semantic space $\mathcal{S}$ generates image $\tilde{X}$ in data space $\mathcal{D}$. The image $\tilde{X}$ is then captioned by $f_{\theta}$ producing a caption $\tilde{Y}$ which can be mapped back to the original latent space as $Y$. The semantic-level error Z=YXZ=Y-X measures the difference between the source semantics $X$ and received semantics $Y$.
@@ 2.1 Formulation
We assume that meaning arises from a set of independent and discrete units called semantic units in a semantic space, and images and captions are observations of some latent variables in this semantic space Peyrard (2019); Zhong (2017). Following Peyrard (2019); Zhong (2017), we define a semantic unit and the semantic space as:
Definition 1
(Semantic Units and Semantic Space) A semantic unit represents an atomic piece of information. The set of all possible semantic units is denoted by $\Omega=\{w_{i}\}_{i=1}^{n}$. A semantic space is an $n$-dimensional space $\mathcal{S}\in[0,1]^{n}$ where the value of the $i$-th dimension of represents the probability of the presence of the corresponding semantic unit $p(w_{i})$.
The $\Omega$ encompasses a wide range of semantic units, similar to how a vocabulary contains diverse words. Each point in $\mathcal{S}$ corresponds to a specific combination of semantic units, corresponding to different meanings. Adopting the semantic space and its probabilistic interpretation in Peyrard (2019); Shao et al. (2022); Niu and Zhang (2024), we can apply the classical information theory Shannon (1948) to operate at the semantic level, and make information in images and captions to be comparable.
@@
Definition 2
[…]
@@ 2.2 Objectives
The image captioning process can be compared to a communication system Shannon (1948) where information source $X$ is converted to signal $\tilde{X}$ by a lossless source encoder $g$, transmitted through a noisy channel $f_{\theta}$, and the received signal $\tilde{Y}$ is losslessly decoded by $h$, giving the final received information $Y$. From this communication system perspective, one might say that the optimal $\theta^{*}$ could just be the one that minimizes the error $Z||$. However, this requirement is unrealistic as it would result in extremely long captions that losslessly encode both high-level semantic information and all the low-level irrelevant information.
[…]
@@
Next, we give a formal definition of the information sufficiency objective of image captioning with importance of semantic units Peyrard (2019).
Definition 3
[…]
@@
In the importance variable, $A_{i}=1$ means the semantic units $w_{i}$ are very important in the image, while $A_{i}=0$ means $w_{i}$ is irrelevant. It behaves similarly to the attention mechanism Bahdanau et al. (2015), which also produces a heatmap between zero to one according to the given input. Note that here the $A$ is not binary, the continual nature of $A$ gives a good property to $J_{\mathrm{suf}}(\theta)$: when the “budget” is limited (as there are also other objectives to optimize), more semantic units with higher importance score will be retained while less important ones will be discarded.
Definition 4
(Minimal Redundancy) The minimal redundancy objective encourages the image captioning model $f_{\theta}$ to eliminate irrelevant information, it is given by measuring the entropy of generated captions:
[…]
@@
Definition 5
(Human Comprehensibility) Let $P_{\tilde{Y}}$ denote the probabilistic distribution of model-generated captions over $\mathcal{D_{\mathrm{caption}}}$, and let $P_{\mathrm{lang}}$ denote the distribution of human interpretable natural language. Given a certain statistical divergence measurement $D$, the human comprehensibility objective is:
@@
3 Proposed Method: PoCa
Figure 2: Global caption tends to only capture the broad overview of the whole image. Local caption alleviates this limitation by providing fine-grained information which can further complement the global caption.
In this section, we introduce the Pyramid of Captions (PoCa) method, which showcases the applicability of our framework to image captionoing research. The key intuition behind the PoCa method is that we can have a more accurate and detailed caption by ensembling multiple captions, as shown in Figure 2. We propose to split an image into multiple local patches generate local captions for each patch, and fuse the captions to obtain a higher-quality caption for the global image.
@@ 3 Proposed Method: PoCa
[…]
@@
We adopt text-only LLMs as the merging function $\sigma_{\mathrm{merge}}$. Table 5 in Appendix provides an example of merging using an LLM, where we instruct it to generate a merged caption that incorporates both local and global information.
@@
4 PoCa Gives Better Captions (Provably)
@@
Assumption 1
(Uncertainty-aware content-dependent error) The error $Z$ produced by the image captioning model $f_{\theta}$ is dependent on the input semantics $X$. Therefore, it can be expressed as a deterministic function of $X$:
$Z_{i}||=\varphi(X_{i}),$ (8)
@@
Assumption 2
(Local-global relationship of image semantics) The $\sigma_{\mathrm{split}}$ splits an image into local patches. The latent semantic variables corresponding to local patches satisfy the following relationship with global semantics:
@@
Assumption 3
(Local-global aggregation of caption semantics) The function $\sigma_{\mathrm{merge}}$ merges the global and global captions. The latent semantic variable corresponding to the merged caption is a weighted combination of the global and local semantics:
@@
Theorem 1
@@
(PoCa method reduces semantic error) Under Assumptions 1-3, the PoCa method is guaranteed to have the smaller error $Z_{\mathrm{merged}}$ than $Z$, i.e.,
[…]
@@
Since $A$ is non-negative, Theorem 1 implies non-decreasing information sufficiency; if merging does not increase redundancy or decrease interpretability, then the overall quality of caption becomes better. This also aligns with the findings in Shi et al. (2024) that smaller-scale models combined can be as effective as a larger-scale model. Additionally, it is worth noting that our assumptions require linear combinations of semantics, while it may not hold in practice for images with more complex structures of semantics.
@@
[…]
$).\displaystyle=(1-\eta)(\varphi(X_{i})-\sum_{j}^{m}\alpha_{j}\varphi(X^{[j]}_{i}$)). (20)
[…]
@@
[…]
@@
5 PoCa Gives Better Captions (Empirically)
5.1 Implementation Details
Image Captioning Models. We employ three groups of Large Vision Language Models (LVLMs) as the image captioning models: LLaVA-1.5 Liu et al. (2023b), MobileVLM v2 Chu et al. (2024), and InternVL Chen et al. (2023b). Among them, LLaVA-1.5 series is a popular LVLM with two variants: LLaVA-1.5-7B and LLaVA-1.5-13B, which adopt Vicuna-7B and Vicuna-13B as their Language Models (LLMs), respectively. MobileVLM v2 is a family of efficient LVLMs with smaller scales, and we utilize its MobileVLM-v2-1.7B and MobileVLM-v2-3B models. InternVL is one of the top-performing publicly available LVLMs. We use its InternVL-Chat-Chinese-V1-2-Plus model, which is based on the Yi-34B LLM and has a total of 40.1B parameters.